OpenStudy (anonymous):

find the zeros: 4x^3+5x^2-3 please help

OpenStudy (anonymous):

Can u write the equation again?, This time using parenthesis. please

OpenStudy (anonymous):

(4x^3)+(5x^2)-3 find the zeros

OpenStudy (anonymous):

give me 5 minuetes.

OpenStudy (anonymous):

np

OpenStudy (anonymous):

dude, do u want just the answer, or do you want a step by step procedure?

OpenStudy (anonymous):

answer im in a rush

OpenStudy (anonymous):

x1= 0.6313819289208555 x2= -0.9406909644604278 + i* 0.5504280704155624 x3= -0.9406909644604278 - i* 0.5504280704155624

OpenStudy (anonymous):

you can find a cubic equation solver at http://www.1728.com/cubic.htm